How do you solve:
Point A (–4,1) is in the standard (x,y) coordinate plane. What must be the coordinates
of point B so that the line x = 2 is the perpendicular bisector of AB?
Thank you

Answer by dlam5(15)

perp. bis. is also a bisector, meaning it must hit in the MIDDLE.
if A(-4,1) then call B(x,y).
if x=2, 2,1 is the midpt. of AB; as midpt. of something is a line; here, it is a horiz. straight line.
Now use midpt. formula.
-4 + x/2 , 1 +y/2 = 2,1
for x; set -4 + x/2=2; find x. x=8
for y; set 1+y/2 = 1; y=1
B(8,1); check.
is 2,1 between A and B? use midpt. to find out...

yes:)


Edd